Three fair dice, each has 6 different faces, are rolled. Let B define the event that no two or no three dice show the same face. What is the probability of B
Answer #1

Total number of outcomes possible = 6*6*6 = 216

Number of outcomes with all different faces = 6*5*4 = 120

Hence,

P(B)

= P(No two or no three die shows the same face)

= P(All faces are different)

= 120/216

= 5/9